OpenStudy (anonymous):

x^2+12x=-36 Use the quadratic formula to solve each equation.

OpenStudy (jhannybean):

Set it to quadratic equation : \(\large \frac{-b \pm \sqrt{b^2 -4ac}}{2a}\)

OpenStudy (anonymous):

This is what I have so far \[x^{2}+12x=-36 \] \[x^{2}+12x+36=-36+36 \] \[x ^{2}+12x +36=0\] \[x=\frac{ -(12)\pm \sqrt{(12)^{2}+4(1)(36)} }{ 2(1) }\] \[x=\frac{ 12\pm \sqrt{144+144} }{ 2 }\] \[x=\frac{ 12\pm \sqrt{288} }{ 2 }\] \[x=\frac{ 12\pm12\sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

whoops but place a - at the first 12 i forgot to place them here on this

OpenStudy (jhannybean):

it should be \[x=\frac{ -(12)\pm \sqrt{(12)^{2}\color{red}-4(1)(36)} }{ 2(1) }\]

OpenStudy (anonymous):

ahhh ok I think thats what i may have done wrong thank you! Im going to check

OpenStudy (anonymous):

ugh, can't believe i missed that! Thank you so much!

OpenStudy (jhannybean):

No problem :) don't forget the -ve infront of the first 12 though!

OpenStudy (anonymous):

it coulde simply x^2+12x=-36 x^2+12x+36=0 (x+6)(x+6)=0 x+6=0 x=-6